OpenStudy (genny7):

Please help! Will give medal! :)))) Calculate the number of moles in 17.8 g of the antacid magnesium hydroxide, Mg(OH)2

OpenStudy (anonymous):

n = m/M n - number of moles m - mass M - molar mass

OpenStudy (anonymous):

Molecular weigh of Mg(OH)2 =12+(8*2)+(1*2)=30 mass= 17.8g if mass=moles * molecular weight then moles =mass/molecular weight moles of Mg(OH)2=17.8/30 = 0.583 to 3d.p

rishavraj (rishavraj):

molar mass of Mg(OH)2 = 24 + (17)*2 = 58. no of moles = m/M m - given mass M- Molar mass